how were the mass % of Cl in each compound calculated?
write the formula.

NaCl-58.4 g/mol - 60.7%Cl
MgCl2- 95.2g.mol - 74.5% Cl
Fecl3- 162.2g/mol - 65.6% Cl
1
Expert's answer
2015-11-03T12:42:25-0500
Mass % was calculated according to the formula w(Cl)=n*m(Cl)/m(MCln), where M=Na, Mg, Fe and n=1-3
For molecules the formula can be rewritten into w(Cl)=M(Cl)/M(MCln), M(Cl)=35.5 g/mol
w(Cl) in NaCl=n*M(Cl)/M(NaCl)=35.5/(35.5+23)=0.607=60.7% here n=1
w(Cl) in MgCl2=n*M(Cl)/M(MgCl2)=2*35.5/(2*35.5+24.3)=0.745=74.5% here n=2
w(Cl) in FeCl3=n*M(Cl)/M(FeCl3)=3*35.5/(3*35.5+56)=0.656=65.6% here n=3